### v2

# Planning of Heuristics: Strategic Planning on Large Language Models with Monte Carlo Tree Search for Automating Heuristic Optimization

Abstract
Heuristics have achieved great success in solving combinatorial optimization problems (COPs). However, heuristics designed by humans require too much domain knowledge and testing time. Given the fact that Large Language Models (LLMs) possess strong capabilities to understand and generate content, and a knowledge base that covers various domains, which offer a novel way to automatically optimize heuristics. Therefore, we propose Planning of Heuristics (PoH), an optimization method that integrates the self-reflection of LLMs with the Monte Carlo Tree Search (MCTS), a well-known planning algorithm. PoH iteratively refines generated heuristics by evaluating their performance and providing improvement suggestions. Our method enables to iteratively evaluate the generated heuristics (states) and improve them based on the improvement suggestions (actions) and evaluation results (rewards), by effectively simulating future states to search for paths with higher rewards. In this paper, we apply PoH to solve the Traveling Salesman Problem (TSP) and the Flow Shop Scheduling Problem (FSSP). The experimental results show that PoH outperforms other hand-crafted heuristics and Automatic Heuristic Design (AHD) by other LLMs-based methods, and achieves the significant improvements and the state-of-the-art performance of our proposed method in automating heuristic optimization with LLMs to solve COPs.
Machine Learning, ICML
1 Introduction
Combinatorial optimization problems (COPs) commonly exist in diverse fields such as national defense(Xing et al., 2024), transportation(Wang & Tang, 2021), industry(Zhao et al., 2024), and communication(Witt et al., 2024). Their substantial theoretical and practical value has made efficient solution of COPs a key research focus in both academia and industry. Heuristics are widely studied to solve COPs and achieve superior performance. The typical heuristics consist of guided local search (GLS) (Voudouris & Tsang, 1999), genetic algorithm (GA) (Holland, 1973), and ant colony optimization (ACO) (Dorigo et al., 1996). Usually, achieving satisfactory solutions with these methods requires human experts to manually adjust the heuristics for each specific problem. Although manually designed heuristics can be effective in many cases, this approach requires much time for experts to design, implement, and validate heuristics. In addition, for many complex COPs, this approach may result in errors. Consequently, Automatic Heuristic Design (AHD) has emerged as a promising alternative (Kumar et al., 2024; Chen et al., 2022). The increasing availability of computational resources further facilitates AHD. By reducing the reliance on specialized domain knowledge, AHD enables us to explore large design spaces, demonstrating substantial potential to address complex COPs.
Given the fact that Large Language Models (LLMs) have demonstrated remarkable capabilities in addressing COPs, due to their broad knowledge understanding and powerful reasoning abilities (Yang et al., 2024). Furthermore, leveraging their extensive training corpora, LLMs benefit from a wider search space compared to traditional evolutionary computation (EC) algorithms, leading to performance improvements (Liu et al., 2024b; Ma, 2024). Recent research has successfully applied LLMs to heuristic generation and evolutionary search processes (Liu et al., 2024a; Ye et al., 2024; Romera-Paredes et al., 2023).
Therefore, in this paper, we introduce a novel AHD approach called Planning of Heuristics (PoH). PoH reframes heuristic optimization as a strategic planning problem to manage the complexity of search spaces. Using strategic planning, PoH iteratively refines heuristics (represented as states) through insightful improvement proposals (actions). Starting with an initial heuristic (state), the system systematically explores the heuristic search space using a tree-based search, prioritizing high-reward trajectories to efficiently navigate this extensive space. Using the Monte Carlo Tree Search (MCTS) planning strategy, PoH can anticipate and simulate future rewards, subsequently using backpropagation to update reward values and guide the search toward more promising trajectories.
Above all, the main contributions of this paper can be summarized as follows.
Figure 1: (a) MCTS planning for heuristic generation. The tree structure enables strategic planning of PoH . (b) A simplified state transition example. The base LLM first initialize (generate) a current heuristic (state), the optimizer LLM gathers improvement suggestions from the task dataset. An optimizer LLM then refines these suggestions, and updates the heuristic (state) accordingly, transitioning to the next state.
  • We propose PoH, a novel automated heuristic optimization method that leverages MCTS planning to strategically and efficiently explore the complex heuristic search space.
  • We show that PoH outperforms several existing automated heuristic design (AHD) methods, including those utilizing LLMs for automatic heuristic optimization.
  • Unlike other frameworks that optimize heuristics with LLMs, PoH is a novel framework that combines iterative self-reflection with planning to optimize heuristics.
2 Related Work
2.1 LLMs for Optimization
LLMs have recently been used to address optimization problems through prompt engineering for specific issues (Yang et al., 2024; Wei et al., 2022). However, relying solely on prompt engineering has proven to have limited effectiveness in complex optimization scenarios. Inspired by the automatic generation of heuristics, researchers have explored combining evolutionary computation (EC) with LLMs to generate and refine heuristics. FunSearch (Romera-Paredes et al., 2023) presents a novel approach that searches within the function space, using LLMs to iteratively improve the quality of generated heuristics within an evolutionary framework. Evolution of Heuristics (EoH) (Liu et al., 2024a) employs natural language to represent heuristic ideas; LLMs first generate natural language descriptions of heuristics, which are then used to produce executable heuristic code. This evolutionary search framework allows for the simultaneous improvement of both the descriptions and the code, contributing to EoH’s effectiveness and efficiency. Similarly, Reflective Evolution (ReEvo) (Ye et al., 2024) enhances the efficiency of heuristic evolution by combining evolutionary search with the self-reflection capabilities of LLMs.
2.2 LLMs with Self-reflection and Planning
Self-reflection is a cognitive process where an individual contemplates their own thoughts, feelings, and actions, enabling the recognition of mistakes during problem-solving and the continuous adjustment of strategies. Similarly, guiding LLMs to engage in self-reflection, allowing them to evaluate their generated content, can effectively improve their problem solving performance (Shinn et al., 2023; Kumar et al., 2024; Huang et al., 2022).
Planning is a crucial tool for agents operating in complex and dynamic environments and making high-quality decisions. Traditional planning methods, which represent problems in a structured format, can leverage efficient search algorithms to generate the correct and optimal solutions (Cheng et al., 2022; Väth et al., 2023; Liu et al., 2023). This motivates research that combines LLMs with planning techniques, such as using MCTS to explore more comprehensive reasoning paths (Wang et al., 2024; Hao et al., 2023).
However, existing LLM-based heuristic optimization methods often rely on evolutionary frameworks or incorporate reflection mechanisms, but typically employ direct iterative algorithms without principled strategies for guided exploration. In contrast, our proposed PoH method synergistically combines self-reflection with planning, leveraging MCTS to efficiently search the vast heuristic space.
3 Methodology
This section introduces PoH, a framework that empowers LLMs to strategically plan a coherent reasoning trajectory to solve a wide range of COPs. We first format the definition of the heuristic optimization problem and then present the proposed PoH framework. Finally, we describe how MCTS planning is employed to effectively explore the vast heuristic space and to identify optimal trajectories.
3.1 Problem Definition
Following a standard setting in heuristic optimization, COP is defined by a solution space $\mathcal{S}$ and an objective function $f:\mathcal{S}\mapsto\mathcal{R}$. Heuristic optimization typically searches within a heuristic space ${H}$ to find an optimal heuristic ${H}^{*}$ that minimizes an evaluation function, formally expressed as ${H}^{*}=\arg\min_{{h}\in{H}}F(h)$. Unlike traditional methods of optimizing heuristics, we leverage LLMs to generate heuristics, enabling exploration of this open heuristic space. Specifically, we evaluate the generated heuristic $H$ on a training set $I$ to maximize performance according to a reward function $\mathcal{R}$, which can be formulated as ${H}^{*}=\arg\max_{{h}\in{H}}\mathcal{R}(p_{\mathcal{B}}({I},{H}))$.
3.2 PoH Framework
Algorithm 1 Planning of Heuristics Method
Inputs:Initial state $s_{0}$, state transition function $p_{\theta}(s^{\prime}|s,a)$, reward function $r_{\theta}(s,a)$, action generator $p_{\alpha}(a|s)$, number of generated actions $h$, depth limit $l$, number of iterations $I$, exploration weight $e$
Initialize: memory of actions $A:\mathcal{S}\mapsto\mathcal{A}$, children $\text{c}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$, rewards $r:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$, State-Action value function $Q:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$, visited count $N:\mathcal{S}\mapsto\mathbb{N}$
1:  for $n\leftarrow 0,\dots,I-1$ do
2:     for $t\leftarrow 0,\dots,l-1$ do
3:        if $A(s_{t})$ is not empty then
4:           $a_{t}\leftarrow\arg\max_{a\in{A}(s_{t})}\left[Q(s_{t},a)+e\cdot\sqrt{\frac{\ln N(s_{t})}{N(\text{c}(s_{t},a))}}\right]$
5:           $s_{t+1}\leftarrow\text{c}(s_{t},a_{t})$, $r_{t}\leftarrow r(s_{t},a_{t})$,
6:           $N(s_{t})\leftarrow N(s_{t})+1$
7:        else
8:           for $j\leftarrow 1,\dots,h$ do
9:              Sample $a_{t}^{j}\sim p_{\alpha}(a|s_{t})$, $s_{t+1}^{j}\sim p_{\theta}(s|s_{t},a_{t}^{j})$,
10:              $r_{t}^{i}\leftarrow r_{\theta}(s_{t},a_{t}^{j})$
11:              Update $A(s_{t})\leftarrow\{a_{t}^{j}\}_{j=1}^{d}$,
12:              $\text{c}(s_{t},a_{t}^{j})\leftarrow s_{t+1}^{j}$,
13:              $r(s_{t},a_{t}^{j})\leftarrow r_{t}^{j}$
14:           end for
15:           $a_{t}\leftarrow\arg\max_{a^{i}_{t}\in{A}(s_{t})}r_{t}^{j}(s_{t},a^{j}_{t})$
16:           $s_{t+1}\leftarrow\text{c}(s_{t},a_{t})$, $r_{t}\leftarrow r(s_{t},a_{t})$,
17:           $N(s_{t})\leftarrow N(s_{t})+1$
18:        end if
19:        if $s_{t+1}$ is an early-stopping state then break
20:     end for
21:     $T^{\prime}\leftarrow$ the actual number of steps
22:     for $t\leftarrow T^{\prime}-1,\dots,0$ do
23:        Update $Q(s_{t},a_{t})$ with $\{r_{t},r_{t+1},\dots,r_{l}\}$
24:     end for
25:  end for
PoH regards the heuristic optimization problem as a Markov Decision Process (MDP), defined by the tuple  $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R})$, with a defined state and action space. Here, $\mathcal{S}$ represents the state space, $\mathcal{A}$ is the action space, $\mathcal{T}$ denotes the transition function, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$, and $\mathcal{R}$ is the reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$. In our work, we formulate the MDP concretely as follows.
State $\mathcal{S}$ is a set of generated heuristics, where $s_{t}\in\mathcal{S}$, and $s_{t}$ is the generated heuristic at step $t$.
Action $\mathcal{A}$ is a set of improvement suggestions, where $a_{t}\in\mathcal{A}$ and $a_{t}$ is the generated improvement suggestion for heuristic $s_{t}$.
Transition function $\mathcal{T}$: When applying an action $a_{t}$ to state $s_{t}$, it yields $s_{t+1}$, which means the heuristic (state $s_{t}$) accepts the improvement suggestion (action $a_{t}$) and updates to a new heuristic (state $s_{t+1}$).
Reward $\mathcal{R}$ is evaluated as $1-\frac{O_{st}-O_{best}}{O_{best}}$ , where $O_{best}$ is known instance-specific baselines, $O_{st}$ is the current heuristic.
As illustrated in Figure 1 (a), given a current state $s_{t}$, PoH iteratively generates an action $a_{t}$ according to $a_{t}\sim p_{\theta}(a|s_{t})$, $p_{{B}}(a|s_{t})$ represents that the base LLM $p_{\theta}$ generates improvement actions $a$ given the current state $s_{t}$. The action generation process, detailed in Figure 1 (b), comprises two steps: first, evaluating the heuristic (state) generated by the base model. This evaluation involves replacing the distance matrix update function of the GLS search algorithm with the generated heuristic and then evaluating it on the training set to obtain the current heuristic’s optimal gap (%). Subsequently, for action generation, the optimizer LLM $p_{\mathcal{B}}$ generates improvement suggestions $a_{t}$ based on the current heuristic $s_{t}$ and the optimal gap (%). PoH then determines the subsequent state using the transition function $p_{\mathcal{\theta}}(s_{t+1}|s_{t},a_{t})$ to update the heuristic. Given the current improvement suggestions (action), the optimizer LLM $p_{\mathcal{\theta}}$ generates a new heuristic (state) that incorporates relevant domain knowledge and effectively addresses model suggestions, which is similar to how AHD via Hyper-Heuristics updates algorithms based on improvement suggestions.
The quality of applying the action $a_{t}$ to state $s_{t}$ is then evaluated by the reward function $r_{t}=r(s_{t},a_{t})$. The reward function is defined as 1 minus the percentage gap between the solution and the optimal solution for an instance, aiming to maximize solution quality and minimize the gap with the optimal solution.
3.3 Planning with Monte Carlo Tree Search
PoH incorporates MCTS as its strategic planning method. A typical MCTS loop consists of the following 4 steps.
Selection Starting from the root node in a search tree, the selection phase traverses the child nodes according to the Upper Confidence Bound (UCT) formula, which balances exploitation and exploration, until a leaf node is reached. The UCT formula is as follows:
$a^{*}=\arg\max_{a\in A(s)}\left[Q(s,a)+e\sqrt{\frac{\ln N(s)}{N(c(s,a))}}\right]$ (1)
Here, ${A}(s)$ represents the set of actions available at node $s$, $N(s)$ represents the number of times node $s$ has been visited, $\text{c}(s,a)$ represents the child node resulting from applying action $a$ to node $s$, and $e$ is a constant that adjusts the degree of exploration. The first term of the formula, $Q(s,a)$, reflects exploitation, while the second term reflects exploration, quantifying the uncertainty associated with the nodes visited less frequently. Specifically, if a node and its child node have been explored insufficiently, the value of the second term will be higher.
Expansion When a leaf node is reached and a terminal state has not yet been achieved, the expansion phase creates one or more new nodes. These new nodes represent future states that may be reached from the current state. During the expansion phase, no evaluation or simulation of these newly added child nodes takes place. Expansion merely increases the structure of the tree in preparation for subsequent simulation steps.
Simulation Starting from the current node, the simulation phase selects actions from all possible actions according to a policy, resulting in state transitions. If a terminal state is not reached, actions continue to be selected until a terminal state is reached.
Back-propagation Back-propagation propagates the simulation result back through the search tree to update node information. When the simulation reaches a terminal state, the $Q$ value of the current node is updated. By continuously updating node information, MCTS guides the search process in more promising directions.
$Q(s_{t},a_{t})=\max_{s_{t},a_{t},r_{t},\dots,s_{l},a_{l},r_{l},s_{l+1}}\operatorname{avg}(r_{t},\dots,r_{l}).$ (2)
Above all, the pseudo code of the PoH method is given as Algorithm 1. The terminal state is when it reaches the predefined maximum depth or meets the early-stopping condition (Line 18 in Algorithm 1). The early stopping condition is activated when the state’s reward is below a minimum threshold or above a maximum threshold. Specifically, the minimum threshold is the average of the rewards of the parent node and the root node, while the maximum threshold is the maximum value among all current nodes. PoH iteratively executes these four phases to explore the heuristic space. The algorithm terminates when a predefined number of iterations is reached, and the best trajectory and corresponding node are then selected for evaluation.
4 Experiments
4.1 Experiments Settings
Tasks and Datasets We evaluated PoH on two classical combinatorial optimization problems: The TSP and the FSSP. The TSP seeks the shortest route that starts at a depot, visits all customer locations exactly once, and returns to the depot. As a canonical combinatorial optimization problem, it serves as a standard benchmark for heuristic methods. Following the setup in (Liu et al., 2024a), we performed heuristic optimization on a training set of 64 TSP instances, each with 100 city nodes randomly distributed in $[0,1]^{2}$ (Kool et al., 2019). We used the average optimal gap, calculated with respect to the solutions generated by Concorde (Applegate et al., 2006), as the evaluation function.
FSSP is a classical problem in production scheduling, which involves scheduling of $n$ jobs on $m$ machines. Each job consists of $m$ operations that must be processed on the machines in the same predetermined order. Each job may have different processing times on each machine. A job can only begin processing on the next machine after completing processing on the previous machine. Each machine can process only one job at a time, and each job can be processed on only one machine at a time. The objective is to minimize the makespan, the total time of finishing all jobs. For heuristic optimization, we used a training set of 64 randomly generated FSSP instances, each with 50 jobs and a varying number of machines (from 2 to 20), consistent with (Liu et al., 2024a). The processing times of the jobs were randomly generated from a uniform distribution between 0 and 1 (Pan et al., 2021). The average optimal makespan, calculated with respect to the solutions generated by (Liu et al., 2024a), served as the evaluation function.
Baselines We compare PoH with three categories of baselines: Guided Local Search (GLS) algorithms, LLM-generated heuristics, and other algorithms for the TSP. GLS algorithms include Knowledge-Guided Local Search (KGLS) (Arnold & Sörensen, 2019a), GNNGLS (Hudson et al., 2022), NeuralGLS (Sui et al., 2023), and a state-of-the-art neural combinatorial optimization (NCO) method (Luo et al., 2023). LLM-generated heuristics include EoH (Liu et al., 2024a) and ReEvo(Ye et al., 2024). Other algorithms include Google Or-Tools (Perron & Furnon, 2023), the Attention Model (AM) (Kool et al., 2019), POMO (Kwon et al., 2021), and LEHD (Luo et al., 2023).
For FSSP, we compared PoH with NEH (Nawaz et al., 1983), NEHFF (Fernandez-Viagas & Framinan, 2014), Local Search (LS), Iterated Local Search (ILS) (Sttzle, 1998), and the AHD methods PFSPNet and PFSPNet_NEH (Pan et al., 2021). NEH and NEHFF are widely recognized as efficient heuristics for this problem. LS and ILS are classical search methods for the FSSP; we used the same operators in LS as those used in PoH.
Implementation details We applied PoH to automatically design Guided Local Search (GLS) heuristics for both the TSP and the FSSP. GLS aims to guide local search away from local optima. A key challenge is to update the objective function to direct the search toward more promising regions. Because the search space is primarily defined by the distance matrix (for the TSP) or the time matrix (for the FSSP), PoH aims to discover heuristics that efficiently update these matrices.
Specifically, for the TSP, following (Arnold & Sörensen, 2019b), the input to the generated heuristic is a distance matrix, and the output is an updated distance matrix. GLS combines a perturbation phase (where edges with higher heuristic values are preferentially penalized) with a local search operator applied to the updated search space.
For the FSSP, we adopted the same GLS framework used for the TSP, employing PoH to generate a heuristic that updates the execution time matrix and determines the jobs to be perturbed. We used the Swap and Relocate local search operators. The inputs to this heuristic are the number of jobs and machines, the current job sequence, and the time matrix; the outputs are the updated job perturbation priority and the updated time matrix.
In our experiments, we used GPT-4.0 as both the base and optimizer models, with temperatures of 0.0 and 1.0, respectively. The MCTS parameters were set as follows: 10 iterations, an expansion width of 5, a maximum depth of 5, and an exploration weight of 2.5.
4.2 Result and Analysis
Traveling Salesman Problem We firstly evaluated the performance of PoH and several other well-studied methods on common instances from TSPLIB (Reinelt, 1991). Table 1 presents the relative distance (%) to the best-known solutions for these methods in a typical subset (selected by (Liu et al., 2024a)) of instances. The complete comparison results with additional TSPLIB instances and other algorithms are provided in the Appendix. As shown in Table 1, EoH achieved the best results on several instances, finding the optimal solution for pr124, kroA150 and u159. In addition, PoH consistently outperformed all other algorithms on all tested instances, demonstrating exceptional robustness and effectiveness, particularly on the larger instance kroB200, where it consistently produced near-optimal solutions.
Table 1: Traveling salesman problem results. Comparison of the relative distance (%) to the best-known solutions (lower is better) for various routing heuristics on a subset of TSPLib instances.
Method rd100 pr124 bier127 kroA150 u159 kroB200
Or-Tools 0.01 0.55 0.66 0.02 1.75 2.57
AM 3.41 3.68 5.91 3.78 7.55 7.11
POMO 0.01 0.60 13.72 0.70 0.95 1.58
LEHD 0.01 1.11 4.76 1.40 1.13 0.64
GNNGLS 0.46 0.76 1.95 2.98 1.02 1.53
GLS 0.01 0.60 0.59 1.75 0.74 1.43
KGLS 0.01 0.08 0.42 0.17 0.96 0.89
EoH 0.01 0.00 0.42 0.00 0.00 0.20
PoH (ours) 0.01 0.00 0.01 0.00 0.00 0.01
Figure 2 compares PoH with two other LLM-based heuristic optimization methods on several large-scale TSPLIB instances. The results demonstrate that PoH significantly outperforms the other two methods on all tested instances. Moreover, PoH’s advantage becomes more pronounced as the problem size increases, indicating its superior ability to find near-optimal solutions for larger instances. Although ReEvo exhibits a relatively small difference compared to PoH on instances with fewer than 1000 nodes, the advantage of PoH is obvious on larger instances.
Figure 2: Comparative analysis of LLM-enhanced heuristics on large-scale TSPLIB instances by showing the optimal gap relative to the optimal solution. PoH achieved significantly smallest gaps especially on problems with larger sizes, which indicates the promising potential of PoH to solve large-scale COPs.
Table 2: Evaluation results of different local search (LS) variants. We report optimality gaps and per-instance execution time.
Method Type TSP20 TSP50 TSP100 TSP200
Opt. gap (%) Time (s) Opt. gap (%) Time (s) Opt. gap (%) Time (s) Opt. gap (%) Time (s)
NeuOpt LS+RL 0.000 0.124 0.000 1.32 0.027 2.67 0.403 4.81
GNNGLS GLS+SL 0.000 0.116 0.052 3.83 0.705 6.78 3.522 9.92
NeuralGLS GLS+SL 0.000 10.005 0.003 10.01 0.470 10.02 3.622 10.12
KGLS GLS 0.004 0.001 0.017 0.03 0.002 1.55 0.284 2.52
EoH GLS+LLMs 0.000 0.563 0.000 1.90 0.025 5.87 0.338 17.52
ReEvo GLS+LLMs 0.000 0.001 0.000 0.02 0.000 0.95 0.306 1.70
PoH (ours) GLS+LLMs 0.000 0.001 0.000 0.02 0.000 0.95 0.227 1.58
In addition, we also conducted experiments on the 64-instance dataset generated by ReEvo (Ye et al., 2024) using seed 1234. The results, presented in Table 2, demonstrate the performance achieved by PoH across different sizes of TSP instances. PoH achieves optimal solutions (0.000% optimality gap) on TSP20, TSP50, and TSP100 instances, matching the best-performing baselines on these smaller instances. More importantly, on larger TSP200 instances, PoH achieves a significantly lower optimality gap of 0.227% compared to all other methods, including both GLS-based methods and other LLM-generated heuristics. This result highlights PoH’s superior ability to scale to larger and more complex TSP instances. Furthermore, PoH achieves these impressive results with competitive execution times, as shown in Table 2. Notably, on TSP200, PoH achieves its superior optimality gap with a runtime comparable to ReEvo, and a runtime much shorter than EoH.
Flow shop scheduling problem As shown in Table LABEL:tab:fssp, PoH demonstrates strong overall performance in all Taillard instances, achieving the lowest average relative makespan in four of the six combinations tested (n20m10, n50m10, n100m10 and n100m20). In particular, PoH achieves significant improvements over traditional methods such as NEH, NEHFF, LS, and ILS1, as well as the deep learning-based methods PFSPNet and PFSPNet_NEH, particularly on instances with 100 jobs. For example, on n100m10, PoH achieves a makespan of 0.11%, significantly lower than the next best result of 0.14% achieved by EoH. Although EoH performs competitively on some smaller instances (n20m20 and n50m10), PoH matches or outperforms it in these cases. On the n50m20 instance, PoH achieves a makespan of 0.49%, slightly higher than the 0.19% achieved by EoH and the 0.47% achieved by LS. However, PoH’s consistent strong performance across the other instances, especially the larger ones, highlights its robustness. The substantial difference in performance between PoH and the deep learning methods PFSPNet and PFSPNet_NEH further underscores the effectiveness of the PoH approach. More detailed results and analysis are provided in the Appendix.
Table 3: Flow shop scheduling problem results. The comparison of the average relative makespan (%) of various baselines on FSSP, evaluated on Taillard instances (lower values indicate better performance).
Method n20m10 n20m20 n50m10 n50m20 n100m10 n100m20
NEH 4.05 3.06 3.47 5.48 2.07 3.58
NEHFF 4.15 2.72 3.62 5.10 1.88 3.73
PFSPNet 14.78 14.69 11.95 16.95 8.21 16.47
PFSPNet_NEH 4.04 2.96 3.48 5.05 1.72 3.56
LS 2.77 2.60 3.33 4.67 1.38 3.51
ILS1 0.33 0.29 1.47 2.13 0.77 2.27
EoH 0.30 0.10 0.19 0.60 0.14 0.41
PoH (ours) 0.19 0.22 0.19 0.49 0.11 0.38
4.3 Heuristic Generalization
We compared PoH’s performance using four commonly used LLMs: GPT-4.0, Gemini-1.5-Pro, GLM-4-Plus, and GPT-3.5-turbo to further demonstrate the generalization of our method across different LLMs. We conducted three independent runs for each LLM on the TSP200 test set, using the same experimental setup for all runs. The results are presented in Table 4. Among the LLMs tested, GPT-4.0 achieved the best performance with PoH, achieving a minimum gap of 0.227% relative to the optimal solution in a single run and an average gap of 0.233% across 3 runs.
Table 4: Comparison of PoH with different LLMs. The average gap (%) to the optimal solution on TSP200 instances.
Method LLM Run 1 Run 2 Run 3 Average
Sampling GPT-4.0 0.762 0.790 0.822 0.791
PoH GPT-3.5-turbo 0.252 0.246 0.234 0.244
PoH GLM-4-Plus 0.276 0.254 0.260 0.260
PoH Gemini-1.5-Pro 0.233 0.236 0.239 0.236
PoH GPT-4.0 0.235 0.227 0.238 0.233
4.4 Ablation on Search Strategies
Figure 3: Ablation study on search methods on TSP200. We compare the optimal gap of four different search methods on the TSP200 test set. The horizontal axis represents the different methods, and the vertical axis represents the percentage of the optimal gap. The error bars indicate the standard deviation of the data, reflecting the stability of the results.
We conducted ablation experiments to further investigate the impact of different search algorithms within the PoH framework. Specifically, we compared PoH using Monte Carlo (MC) search, depth-first Greedy Search (Greedy), and beam search. For each search algorithm, we maintain identical state transitions and action generation processes, only substituting MCTS with the respective algorithm. MC randomly samples and selects an action. Greedy selects the action with the highest immediate reward at each step. The beam search retains the most promising paths (determined by the beam width) during action selection and expands along these paths to find a solution. The total number of explored heuristics was kept constant in all search algorithms.
Figure 3 demonstrates that both Greedy and Beam search outperform MC within the PoH framework, indicating PoH’s capacity for iterative improvement through guided exploration. Furthermore, the beam search performs better than Greedy. In particular, MCTS achieves the smallest deviation from the optimal solution and exhibits greater stability (smaller error bars).
Figure 4: Convergence curves on TSP200 with varying iterations. The vertical axis represents the optimal gap to the baseline for TSP200. The horizontal axis represents the number of MCTS iterations. The shaded areas represent the standard deviation of the current node’s performance.
Table 5: Comparison of exploration efficiency with different search strategy. The average gap (%) to the optimal solution on TSP200 instances.
Method Explored heuristics gap(%)
Greedy Search 34 0.530
Beam Search 72 0.260
MCTS (our PoH) 60 0.227
Figure 5: Convergence curves on TSP200 with varying tree depths. The vertical axis represents the optimal gap to the baseline for TSP200. The horizontal axis represents the MCTS tree depth. The shaded areas represent the standard deviation of the current node’s performance.
4.5 Exploration Efficiency Analysis
To demonstrate PoH’s effectiveness in exploring the prompt space through strategic planning, we compared it with several other search algorithms to analyze exploration efficiency. Specifically, we compared PoH with Greedy Search (equivalent to Greedy in ablation study but with a expand width of 1) and beam search (with a beam width of 3). These two search algorithms explored 34 and 72 heuristics, respectively. As shown in Table 5, increasing the beam width from 1 to 3 improves performance, but neither surpasses MCTS (used in our proposed PoH). Furthermore, beam search requires a higher computational cost. These results demonstrate that PoH strategically and efficiently traverses the heuristic space and effectively finds optimal trajectories.
4.6 Convergence Analysis
Figure 6: This figure illustrates the state-action transition process of MCTS within PoH for training on the TSP. An initial heuristic (state), $s_{0}$, is generated by a base model. In each transition, an optimizer model proposes improvement suggestions (actions) based on the current state. PoH then generates a new heuristic (state) considering both the previous state and these suggestions. Evaluated on TSP200 test instances, this process reduced the gap between the obtained solution and the optimal solution from 0.483% to 0.233%.
We conducted a convergence analysis of PoH’s learning process, visualizing the optimal gap with respect to both the number of MCTS iterations and tree depth. We used TSP200 as a representative test instance for clearer visualization. Figure 4 shows the convergence with varying iterations, where the horizontal axis represents the number of iterations and the vertical axis represents the optimal gap. The performance of PoH during training and testing is shown by blue and yellow lines, respectively. Figure 5 shows the convergence with varying tree depths, where the horizontal axis represents the tree depth and the vertical axis represents the optimal gap. In both figures, the optimal gap decreases as the number of iterations and tree depth increase, demonstrating PoH’s ability to iteratively improve the generated heuristic.
4.7 Qualitative Analysis
Figure 6 provides a qualitative analysis of how PoH iteratively refines its heuristic (state) based on improvement suggestions (actions), demonstrating its capacity for strategic planning. The figure depicts four states ($s_{0}$ to $s_{3}$) and three actions ($a_{0}$ to $a_{2}$) within a TSP training trajectory, showcasing the continuous optimization of the heuristic from its initial state ($s_{0}$). Each subsequent state incorporates the suggestions from previous iterations, resulting in a heuristic with a progressively decreasing gap from the optimal solution on the test instance, reducing from 0.483% to 0.233%.
5 Conclusion
This paper proposed Planning of Heuristics (PoH), a novel framework for automated heuristic design that combines Large Language Models (LLMs) with the planning strategies of MCTS. PoH strategically searches within the vast heuristic space through MCTS-based planning algorithm. Furthermore, PoH leverages LLM self-reflection to identify shortcomings in generated heuristics and propose targeted improvements. We evaluated PoH on two benchmark combinatorial optimization problems: the Traveling Salesman Problem (TSP) and the Flow Shop Scheduling Problem (FSSP). Experimental results demonstrate that PoH outperforms both other LLM-optimized heuristics and manually designed heuristics.
For future work, researchers are encouraged to apply our proposed method to solve large-scale COPs in practice. MCTS simulation may cost much time, incorporating cheaper simulation methods such as Gumbel MCTS should be promising on sovling large-scale COPs .
6 Acknowledgment
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.